# Quantum-Informed Machine Learning for Chaotic Systems

### v2

@@ II Applications
The selected test cases span increasing levels of physical and dynamical complexity. We begin with the Kuramoto–Sivashinsky equation, a one-dimensional PDE exhibiting spatiotemporal chaos. Next, we consider two-dimensional Kolmogorov flow, governed by the 2D incompressible Navier–Stokes equations in a periodic domain. Finally, we examine cross-sections of a three-dimensional turbulent channel flow, simulated using the lattice Boltzmann method (LBM), which is numerically equivalent to solving the Navier–Stokes equations [57]. While the full flow field is simulated numerically, the extracted two-dimensional cross-sectional datasets used in our analysis do not satisfy any closed-form governing equation. This distinction highlights a further step in complexity: moving from lower-dimensional systems with explicit dynamics to high-dimensional observational data without accessible evolution laws, thereby testing the capacity of QIML to generalise across different scenarios.
Throughout this study, ‘ground truth’ or ‘raw data’ refers to the high-resolution simulation data generated by established numerical solvers. While this raw simulation data has a large memory footprint, the quantum-informed priors offer substantial memory efficiency. They capture the flow’s essential statistical features using fewer than 300 trainable parameters, occupying only a few megabytes of storage, which corresponds to a data storage reduction of over two orders of magnitude compared to the raw data. This approach enables a substantial reduction in data storage requirements compared to using the raw simulation data (see section Method E and Supplementary Sections S2–S3). The classical components of the models were trained on a single NVIDIA A100 80GB GPU using default settings with 32-bit floating-point precision (for details see Supplementary Section S4.3).
II.1 Kuramoto-Sivashinsky equation
@@
where denotes a scalar field representing the state of the system, with representing the spatial domain and the time; and are constants which are set to 1 for this study. The spatial domain is discretised using equidistant grid points. Raw data are generated as described in Supplementary Information S4.1. The dataset consists of trajectories, where each trajectory is made up of time frames. The spatial domain for each frame is discretised using grid points. We partition the trajectories chronologically into training, validation, and test sets. For this system, both the classical Koopman surrogate (without Q-prior) and the main QIML model (with Q-prior) are trained for epochs. The Q-prior for the QIML configuration is generated by a QCBM using qubits ( basis states) and trainable parameters. Full architectural and training details for both configurations are provided in Supplementary Information S7.
@@
Fig. 2 presents the QIML evaluation results on the Kuramoto–Sivashinsky system, comparing two configurations against the numerical simulation: a classical surrogate model trained without Q-priors (without Q-prior), and a quantum-informed surrogate model incorporating a QCBM prior trained in simulation (with Q-prior). Both surrogate models perform inference in an auto-regressive manner: starting from an initial condition at time step , the model predicts the next frame at each subsequent step and recursively continues this process up to time steps to ensure long-term statistics. Panel (a) shows the evolution alone the time: with 512 initial states and 500 prediction timesteps and the corresponding relative error evolution. The error metric definition is presented in Supplementary S9. From top to bottom, each row corresponds to the ground truth, the ML model without Q-prior configuration, and the QIML model with Q-prior configuration. These sub-panels are computed over the test dataset to visualize long-term statistical characteristics of the predictions. In the “without Q-Prior” configuration, the model reproduces the general structure of the ground truth. However, over time, the predicted stripe patterns drift upward, revealing noticeable discrepancies in the associated error map. This indicates that the model accumulates errors during long-term prediction, likely due to the absence of prior information to guide the dynamics. In contrast, the “with Q-Prior” model demonstrates much closer alignment with the ground truth. The predicted fields preserve fine-scale features and capture the parallel stripe structures effectively. Additionally, the corresponding error maps exhibit substantially lower and more spatially diffuse errors. These results suggest that incorporating the Q-Prior enhances the model’s stability and accuracy in long-term forecasting. Panel (b) presents the probability density function (PDF) of for each configuration. The distribution is obtained via the entire spatiotemporal test domain (512 states in space and 500 prediction steps in time), allowing for a comparison of global statistical fidelity. The raw test data (solid blue line) serves as the ground truth. The model without Q-Prior (green dashed line) deviates from the raw data, particularly in the tails of the distribution, where it tends to overestimate the probability. In contrast, the model with Q-Prior (orange dash-dotted line) closely matches the ground truth across the full velocity range, especially in the tails. This demonstrates that incorporating the Q-Prior generated by the quantum emulator allows the surrogate model to more accurately reproduce the statistical characteristics of the system, including rare events. Panel (c) shows the energy spectrum as a function of the wave number , comparing the raw data, the surrogate model with and without Q-Prior. The energy spectrum of the raw data follows a smooth decay with increasing , reflecting the distribution of energy across spatial frequencies. The model without Q-Prior overestimates the energy in the large-scale, deviating from the ground truth. However, the model with Q-Prior closely aligns with the raw data across all wave numbers, able to capture both large scale and small-scale (high-) behaviour of the system. To further visualize the invariant structures underlying the KS system, we project both the ground truth and the Q-Prior configuration into the phase space, as shown in Panel (d). The resulting density plots depict the support of the system’s invariant measure, providing insight into its long-term statistical behaviour. Over extended prediction, the model with Q-Prior closely aligns with the ground truth in phase space, accurately capturing the geometry and concentration of the invariant set. This qualitative agreement highlights the Q-Prior model’s ability to preserve the system’s underlying statistical structures over time. Panel (e) illustrates the temporal autocorrelation of the velocity field, averaged over all spatial points, as a function of the dimensionless time lag . Autocorrelation measures how strongly the system retains correlation with its past states over time. Specifically, its calculation is presented in Supplementary S9. In this plot, all cases exhibit a decay in autocorrelation, eventually showing no correlation with the initial state. The configuration with Q-Prior demonstrates stronger short-term correlation (the same as the ground truth data) compared to the model without Q-Prior, indicating better preservation of temporal dynamics in the early prediction period. These outcomes demonstrate that our designed Q-prior can regularise the classical ML model, preserving both long-scale and small-scale structures, and improving long-term statistical accuracy.
II.2 2D Kolmogorov flow
@@
where is the velocity field, represents the pressure. Here denotes the velocity field on the 2D cases. In the algorithmic expressions, we revert to the simpler symbol , using it as a generic placeholder for both 1D scalar and 2D velocities, since the learning procedure applies uniformly across dimensionalities. is the kinematic viscosity, and the forcing term is given by for some amplitude and wave number . The domain is periodic in both and directions: with grids. The data source is presented in Supplementary Information S4.2. We use for training, for validation, and for testing. For this system, both the classical Koopman surrogate (without Q-prior) and the main QIML model (with Q-prior) are trained for epochs. For QIML, the QCBM uses qubits ( basis states) and trainable parameters. Detailed network architectures, training hyperparameters, and data generation procedures are provided in Supplementary Information S7.
@@
Figure 3 reports the comparative performance of the baseline ML model (without Q-prior) and the quantum-informed ML model (with Q-prior) on the two-dimensional Kolmogorov flow dataset. Panel (a) shows the time-averaged velocity field and the corresponding turbulent kinetic energy (TKE) field across three configurations: the ground truth from simulation, the classical model trained without Q-prior, and the quantum-informed model trained with Q-prior. All models are evaluated on the same test data. Both the “with Q-Prior” and “without Q-Prior” models perform auto-regressive (one frame in one frame out) inference over a horizon of time steps. The visualisations span a two-dimensional domain of size in the – plane. The time-averaged velocity field is plotted on the left column and the corresponding TKE field is shown on the right column. The ground truth velocity field (top left) exhibits coherent, stripe-like structures, and its TKE field (top right) reveals concentrated regions of high energy, indicating the presence of organized flow features. The model without Q-Prior (middle row) shows degraded spatial coherence in the velocity field, with disrupted patterns and more noise-like structures. In contrast, the model with Q-Prior (bottom row) recovers a more organized velocity field, better preserving the stripe patterns observed in the ground truth. Its associated TKE field also aligns more closely with the ground truth. Panel (b) presents the ensemble-averaged energy spectrum as a function of the wavenumber for the raw data (solid blue line), the surrogate model without Q-Prior (green dashed line), and the model with Q-Prior (orange dash-dotted line). The spectrum from the model without Q-prior overestimates the energy spectrum at low and intermediate wave numbers. However, the Q-prior aligns closer to the raw data across the whole spectrum. Then, we further examine the overall velocity distributions. As illustrated in Panel (c), we compare the predicted velocity probability density functions from both surrogate models. The raw data exhibit a sharply peaked distribution with heavy tails, characteristic of the system’s inherent velocity fluctuations. Both surrogate models capture the general symmetric shape of the distribution but tend to slightly underestimate the peak and overestimate the spread. Among them, the model with Q-Prior more closely matches the raw data, particularly around the peak and in the tails, indicating improved accuracy in reproducing the system’s statistical behaviour. These results further support the conclusion that incorporating the Q-Prior enhances the surrogate model’s ability to recover velocity statistics. We also examined the auto-correlation of the field along time. Panel (d) plots the autocorrelation of the velocity field as a function of dimensionless time . The correlation is computed over 100 randomly sampled locations, with shaded regions indicating one standard deviation. Both surrogates reproduce long-term decay in autocorrelation followed by fluctuations around zero, consistent with the expected loss of memory in a chaotic or turbulent system. To further analyse the model, Panel (e) reports the normalised relative auto-correlation error with respect to the ground truth. The raw data serves as the reference baseline with zero error. The model without Q-Prior shows a gradual decrease in relative error over time but maintains a consistently higher error floor. In contrast, the model with Q-Prior exhibits a more rapid reduction in normalized relative error and achieves lower relative error throughout the prediction horizon.
@@
These observations confirm that the quantum-informed Q-prior can supply additional information that improves classical ML fidelity and velocity statistics, and enhances long-horizon accuracy.
II.3 Turbulent Channel Flow Inflow Generation
@@
We further evaluate our QIML framework on turbulent channel flow, a representative case of wall-bounded turbulence. In this setting, the training data is obtained from the mid-plane cross-section of a fully developed periodic turbulent channel flow at friction Reynolds number . Note that the reference data has been validated with the DNS data[58] (See supplementary information S5.5). The reference data is generated using the LBM solver, which serves as an efficient alternative to directly solving the Navier-Stokes equations[57, 59]. A large eddy simulation (LES) formulation is employed to capture the large-scale turbulent structures [60, 24]. The governing equations used for generating the data are provided in the Supplementary Information S.5. The detailed description of the data generation setup is described in Supplementary Information S.8.3. The channel flow snapshots are split into training, validation and test sets. For this system, both the classical Koopman surrogate (without Q-prior) and the main QIML model (with Q-prior) are trained for epochs. For QIML, we employ a -qubit QCBM ( basis states) with trainable parameters. Detailed network architectures, training hyperparameters, and data generation procedures are provided in Supplementary Information S7.
@@
Fig. 4 presents the evaluation of the quantum-informed machine learning framework on the turbulent channel flow dataset. Our ML models perform auto-regressive (one frame in one frame out) inference over a horizon of time steps. The visualisations span a two-dimensional domain of size in the – plane. Panel (a) shows a comparison of time-averaged velocity magnitudes across a domain for the ground truth, the surrogate model without Q-Prior, and two Q-Prior-enhanced variants: the emulator and real superconducting IQM device. The colourbar on the right indicates the magnitude of , with higher values in yellow and lower values in blue. The ground truth (top left) displays smooth large-scale variations in the velocity field. The model without Q-Prior (top right) introduces noticeable small-scale noise and fails to capture the smooth gradients seen in the true field. In contrast, both Q-Prior (with emulator and with IQM device) models (bottom row) closely match the ground truth, preserving the dominant spatial structures and exhibiting lower spurious variability. Panel (b) plots the ensemble energy spectrum as a function of the wavenumber for the raw data (solid blue line), the surrogate model without Q-Prior (green dashed line), the model with Q-Prior (orange dash-dotted line), and the model with Q-Prior IQM (purple dotted line). The model without Q-Prior deviates from the ground truth at higher wavenumbers, with discrepancies arising at smaller spatial scales. In contrast, both Q-Prior variants better capture the energy distribution across all scales. Panel (c) compares the normalised TKE fields across a spatial domain for the ground truth, the model without Q-Prior, and two Q-Prior-based configurations: the Emulator and IQM. The ground truth (top left) displays spatially structured regions of elevated TKE, particularly near the domain boundaries, reflecting characteristic flow features. The model without Q-Prior (top right) fails to capture any meaningful TKE patterns, instead producing a near-zero field, which indicates an almost static prediction over time. In contrast, both Q-Prior-based models (bottom row) recover structured TKE fields that closely resemble the ground truth. These models preserve the dominant energy distribution and capture the spatial organization of turbulent structures more accurately. Panel (d) shows the dimensionless velocity profile as a function of the dimensionless wall-normal coordinate . Results are compared with the raw data, log-law and DNS reference. All surrogates reproduce the canonical log-law behaviour and match well with the reference data, DNS data and log law.

### v3

@@ III.1 Numerical set up
The selected test cases span increasing levels of physical and dynamical complexity. We begin with the Kuramoto–Sivashinsky (KS) equation, a PDE exhibiting spatiotemporal chaos. The solution is discretised into 512 spatial grid points, resulting in a 512-dimensional state vector at each time step. Next, we consider two-dimensional Kolmogorov flow on a spatial grid (dimension ), governed by the 2D incompressible Navier–Stokes equations in a periodic domain, where the quantum generator is trained in emulation to extract turbulent invariant measures. Finally, we examine cross-sections of a three-dimensional turbulent channel flow (TCF), simulated using the lattice Boltzmann method (LBM), which is numerically equivalent to solving the Navier–Stokes equations [74]. For the training process, the raw cross-sections are down-sampled and discretised onto a spatial grid, resulting in a 4096-dimensional state vector. The Q-Prior for TCF is trained on superconducting quantum processors. While the full flow field is simulated numerically, the extracted two-dimensional cross-sectional datasets used in our analysis do not satisfy any closed-form governing equation. This distinction highlights a further step in complexity: moving from lower-dimensional systems with explicit dynamics to high-dimensional observational data without accessible evolution laws, thereby testing the capacity of QIML to generalize across different scenarios. Throughout this study, ‘ground truth’ or ‘raw data’ refers to the high-resolution simulation data generated by established numerical solvers. While this raw simulation data has a large memory footprint, the quantum-informed priors offer parameter efficiency and memory advantage. They capture the flow’s essential statistical features using fewer than 300 trainable parameters, occupying only a few megabytes of storage, which corresponds to a data storage reduction of over two orders of magnitude compared to the raw data. This approach enables a substantial reduction in data storage requirements compared to using the raw simulation data (see section Method E and Appendix B–C). The classical components of the models were trained on a single NVIDIA A100 80GB GPU using default settings with 32-bit floating-point precision (for details, see Appendix D).
III.2 Kuramoto-Sivashinsky equation
@@
where denotes a scalar field representing the state of the system, with representing the spatial domain and the time; and are constants which are set to 1 for this study. The spatial domain is discretised using equidistant grid points. Raw data are generated as described in Appendix D. The dataset consists of trajectories, where each trajectory is made up of time frames. The spatial domain for each frame is discretised using grid points. We partition the trajectories chronologically into training, validation, and test sets. For this system, both the classical Koopman machine learning (without Q-Prior) and the main QIML model (with Q-Prior) are trained for epochs. The Q-Prior for the QIML configuration is generated by a sample-based quantum generator using qubits ( basis states) and trainable parameters. Full architectural and training details for both configurations are provided in Appendix G.
@@
Fig. 2 presents the QIML evaluation results on the Kuramoto–Sivashinsky system, comparing two configurations against the numerical simulation: a classical machine learning model trained without Q-Priors (without Q-Prior), and a quantum-informed machine learning model incorporating a prior trained in simulation (with Q-Prior). Both machine learning models perform inference in an auto-regressive manner: starting from an initial condition at time step , the model predicts the next frame at each subsequent step and recursively continues this process up to time steps to ensure long-term statistics. Panel (a) shows the evolution alone the time: with 512 initial states and 500 prediction timesteps, and the corresponding relative error evolution. The error metric definition is presented in Appendix I. From top to bottom, each row corresponds to the ground truth, the ML model without Q-Prior configuration, and the QIML model with Q-Prior configuration. These sub-panels are computed over the test dataset to visualize long-term statistical characteristics of the predictions. In the “without Q-Prior” configuration, the model reproduces the general structure of the ground truth. However, over time, the predicted stripe patterns drift upward, revealing noticeable discrepancies in the associated error map. This indicates that the model accumulates errors during long-term prediction, likely due to the absence of prior information to guide the dynamics. In contrast, the “with Q-Prior” model demonstrates much closer alignment with the ground truth. The predicted fields preserve fine-scale features and capture the parallel stripe structures effectively. Additionally, the corresponding error maps exhibit substantially lower and more spatially diffuse errors. These results suggest that incorporating the Q-Prior enhances the model’s stability and accuracy in long-term forecasting. Panel (b) presents the probability density function (PDF) of for each configuration. The distribution is obtained via the entire spatiotemporal test domain (512 states in space and 500 prediction steps in time), allowing for a comparison of global statistical fidelity. The raw test data (solid blue line) serves as the ground truth. The model without Q-Prior (green dashed line) deviates from the raw data, particularly in the tails of the distribution, where it tends to overestimate the probability. In contrast, the model with Q-Prior (orange dash-dotted line) closely matches the ground truth across the full velocity range, especially in the tails. This improved statistical fidelity is confirmed by a quantitative analysis of the overall prediction error; the inclusion of the Q-Prior results in a 17.25% reduction in the Mean Squared Error (MSE) of the predicted fields with respect to the ground truth over the first 100 steps. This effect becomes even more pronounced in the distribution tails (events where ), where the error is suppressed by nearly two orders of magnitude. Moreover, an analysis of the spectrum demonstrates superior performance across all energy scales, achieving a 29.36% reduction in MSE for this metric. This demonstrates that incorporating the Q-Prior generated allows the machine learning model to more accurately reproduce the statistical characteristics of the system, including rare events. Panel (c) shows the energy spectrum as a function of the wave number , comparing the raw data, the machine learning model with and without Q-Prior. The energy spectrum of the raw data follows a smooth decay with increasing , reflecting the distribution of energy across spatial frequencies. The model without Q-Prior overestimates the energy in the large-scale, deviating from the ground truth. However, the model with Q-Prior closely aligns with the raw data across all wave numbers, able to capture both large-scale and small-scale (high-) behavior of the system. To further visualize the invariant structures underlying the KS system, we project both the ground truth and the Q-Prior configuration into the phase space, as shown in Panel (d). The resulting density plots depict the support of the system’s invariant measure, providing insight into its long-term statistical behavior. Over an extended prediction, the model with Q-Prior closely aligns with the ground truth in phase space, accurately capturing the geometry and concentration of the invariant set. This qualitative agreement highlights the Q-Prior model’s ability to preserve the system’s underlying statistical structures over time. Panel (e) illustrates the temporal autocorrelation of the velocity field, averaged over all spatial points, as a function of the dimensionless time lag . Autocorrelation measures how strongly the system retains correlation with its past states over time. Specifically, its calculation is presented in Appendix I. In this plot, all cases exhibit a decay in autocorrelation, eventually showing no correlation with the initial state. The configuration with Q-Prior demonstrates stronger short-term correlation (the same as the ground truth data) compared to the model without Q-Prior, indicating better preservation of temporal dynamics in the early prediction period. These outcomes demonstrate that our designed Q-Prior can regularize the classical ML model, preserving both long-scale and small-scale structures, and improving long-term statistical accuracy.
III.3 2D Kolmogorov flow
@@
where is the velocity field, represents the pressure. Here denotes the velocity field on the 2D cases. In the algorithmic expressions, we revert to the simpler symbol , using it as a generic placeholder for both 1D scalar and 2D velocities, since the learning procedure applies uniformly across dimensionalities. is the kinematic viscosity, and the forcing term is given by for some amplitude and wave number . The domain is periodic in both and directions: with grids. The data source is presented in Appendix D. We use for training, for validation, and for testing. For this system, both the classical Koopman machine learning (without Q-Prior) and the main QIML model (with Q-Prior) are trained for epochs. For QIML, the quantum generator uses qubits ( basis states) and trainable parameters. Detailed network architectures, training hyperparameters, and data generation procedures are provided in Appendix G.
@@
Fig. 3 presents the predicted streamwise velocity fields over a 20-step autoregressive rollout (), where the time is normalized as . The snapshots qualitatively demonstrate that the QIML model preserves the coherent vortical structures of the ground truth with higher fidelity and for a longer duration compared to its classical counterpart, in which these features decay more rapidly into a disorganized state. Fig. 4 reports the comparative performance of the baseline ML model (without Q-Prior) and the quantum-informed ML model (with Q-Prior) on the two-dimensional Kolmogorov flow dataset. Panel (a) shows the time-averaged velocity field and the corresponding turbulent kinetic energy (TKE) field across three configurations: the ground truth from simulation, the classical model trained without Q-Prior, and the quantum-informed model trained with Q-Prior. All models are evaluated on the same test data. Both the “with Q-Prior” and “without Q-Prior” models perform auto-regressive (one frame in one frame out) inference over a horizon of time steps. The visualizations span a two-dimensional domain of size in the – plane. The time-averaged velocity field is plotted in the left column, and the corresponding TKE field is shown in the right column. The ground truth velocity field (top left) exhibits coherent, stripe-like structures, and its TKE field (top right) reveals concentrated regions of high energy, indicating the presence of organized flow features. The model without Q-Prior (middle row) shows degraded spatial coherence in the velocity field, with disrupted patterns and more noise-like structures. In contrast, the model with Q-Prior (bottom row) recovers a more organized velocity field, better preserving the stripe patterns observed in the ground truth. Its associated TKE field also aligns more closely with the ground truth. Panel (b) presents the ensemble-averaged energy spectrum as a function of the wavenumber for the raw data (solid blue line), the machine learning model without Q-Prior (green dashed line), and the model with Q-Prior (orange dash-dotted line). The spectrum from the model without Q-Prior overestimates the energy spectrum at low and intermediate wave numbers. However, the Q-Prior aligns closer to the raw data across the whole spectrum. Then, we further examine the overall velocity distributions. As illustrated in panel (c), we compare the predicted velocity probability density functions from both machine learning models. The raw data exhibit a sharply peaked distribution with heavy tails, characteristic of the system’s inherent velocity fluctuations. Both machine learning models capture the general symmetric shape of the distribution, but tend to slightly underestimate the peak and overestimate the spread. Among them, the model with Q-Prior more closely matches the raw data, particularly around the peak and in the tails, indicating improved accuracy in reproducing the system’s statistical behavior. This is quantitatively supported by a 6.57% reduction in the MSE for the Q-Prior model’s predictions compared to the classical model without the prior over the first 100 steps. This advantage is particularly pronounced in the high-probability peak of the distribution (where ), where the MSE reduction for the Q-prior model increases to 10.39%. On a global scale, this superior performance is also reflected in the energy spectrum, where the Q-Prior achieves a 14.16% MSE reduction. These results further support the conclusion that incorporating the Q-Prior enhances the machine learning model’s ability to recover velocity statistics. We also examined the autocorrelation of the field over time. Panel (d) plots the autocorrelation of the velocity field as a function of dimensionless time . The correlation is computed over 100 randomly sampled locations, with shaded regions indicating one standard deviation. Both models reproduce long-term decay in autocorrelation followed by fluctuations around zero, consistent with the expected loss of memory in a chaotic or turbulent system. To further analyze the model, panel (e) reports the normalized relative auto-correlation error with respect to the ground truth. The raw data serves as the reference baseline with zero error. The model without Q-Prior shows a gradual decrease in relative error over time but maintains a consistently higher error floor. In contrast, the model with Q-Prior exhibits a more rapid reduction in normalized relative error and achieves lower relative error throughout the prediction horizon.
@@
These observations confirm that the quantum-informed Q-Prior can provide additional information that improves classical ML fidelity and velocity statistics, and enhances long-horizon accuracy.
III.4 Turbulent channel flow inflow generation
@@
We further evaluate our QIML framework on turbulent channel flow, a representative case of wall-bounded turbulence. In this setting, the training data is obtained from the mid-plane cross-section of a fully developed periodic turbulent channel flow at friction Reynolds number . Note that the reference data has been validated with the DNS data [75] (See Appendix E). The reference data is generated using the LBM solver, which serves as an efficient alternative to directly solving the Navier-Stokes equations [74, 76]. A large eddy simulation (LES) formulation is employed to capture the large-scale turbulent structures [77, 24]. Figure 5 a shows the instantaneous velocity field extracted from the high-fidelity turbulent channel flow simulation, with the red ellipse marking the cross-section used as the source data for inflow synthesis. The governing equations used for generating the data are provided in Appendix E. The detailed description of the data generation setup is in Appendix H. The channel flow snapshots are split into training, validation, and test sets. For this system, both the classical Koopman machine learning model (without Q-Prior) and the main QIML model (with Q-Prior) are trained for epochs. For QIML, we employ a -qubit quantum generator ( basis states) with trainable parameters. Detailed network architectures, training hyperparameters, and data generation procedures are provided in Appendix G. Figure 5 b depicts the corresponding synthetic turbulent inflow generated by the QIML model, demonstrating its capability to produce physically consistent inflow conditions for large-eddy simulations [77].
@@
Figure 6 presents the evaluation of the quantum-informed machine learning framework on the turbulent channel flow dataset. Our ML models perform auto-regressive (one frame in one frame out) inference over a horizon of time steps. The visualizations span a two-dimensional domain of size in the – plane. Panel (a) shows a comparison of time-averaged velocity magnitudes across a domain for the ground truth, the machine learning model without Q-Prior, and two Q-Prior-enhanced variants: the emulator and real superconducting IQM device. The colorbar on the right indicates the magnitude of , with higher values in yellow and lower values in blue. The ground truth (top left) displays smooth, large-scale variations in the velocity field. The model without Q-Prior (top right) introduces noticeable small-scale noise and fails to capture the smooth gradients seen in the true field. In contrast, both Q-Prior (with emulator and with IQM device) models (bottom row) closely match the ground truth, preserving the dominant spatial structures and exhibiting lower spurious variability. Panel (b) plots the ensemble energy spectrum as a function of the wavenumber for the raw data (solid blue line), the machine learning model without Q-Prior (green dashed line), the model with Q-Prior (orange dash-dotted line), and the model with Q-Prior IQM (purple dotted line). The model without Q-Prior deviates from the ground truth at higher wavenumbers, with discrepancies arising at smaller spatial scales, which indicating a failure to capture the essential small-scale physical dynamics of the system. As we will demonstrate in the following section IV, this is a common failure mode for state-of-the-art classical architectures. In contrast, both Q-Prior variants better capture the energy distribution across all scales. Panel (c) compares the normalized TKE fields across a spatial domain for the ground truth, the model without Q-Prior, and two Q-Prior-based configurations: the emulator and IQM. The ground truth (top left) displays spatially structured regions of elevated TKE, particularly near the domain boundaries, reflecting characteristic flow features. The model without Q-Prior (top right) fails to capture any meaningful TKE patterns, instead producing a near-zero field, which indicates an almost static prediction over time. In contrast, both Q-Prior-based models (bottom row) recover structured TKE fields that closely resemble the ground truth. These models preserve the dominant energy distribution and capture the spatial organization of turbulent structures more accurately. Panel (d) shows the dimensionless velocity profile as a function of the dimensionless wall-normal coordinate . Our results are compared with the raw data, log-law and DNS reference. All machine learning models reproduce the canonical log-law behavior and match well with the reference data, DNS data and log law.